OpenStudy (anonymous):

suppose that ln 2 = a and ln 5 = b. use properties of logarithm in terms of a and b. ln \[\sqrt[7]{20}\]

OpenStudy (anonymous):

\[\ln(\sqrt[7]{20}) = \ln(20^{\frac{1}{7}}) = \frac{1}{7}\ln(20) = \frac{1}{7}\ln(2^2*5) = \frac{1}{7}(2\ln(2)+\ln(5)) = \frac{1}{7}(2a+b)\]

OpenStudy (anonymous):

thank you.